# Recursively-Constrained Partially Observable Markov Decision Processes

### v2

@@ 5 Dynamic Programming for RC-POMDPs
where $\alpha_{c}$ refers to a cost $\alpha$-vector of $\pi_{c_{min}}$. This can be solved efficiently with the maximin LP (Williams 1990):
$\max_{t,b}\,t\;\;\;\text{ s.t. }\;\;\;\alpha_{c}^{T}b\geq t,\;\;(\alpha_{r},\alpha_{c})\in\Gamma,\;\;b\in\Delta(S).$ (21)
Sampling
ARCS uses a mixture of random search (Spaan and Vlassis 2005) and heuristic search (Smith and Simmons 2005; Kurniawati, Hsu, and Lee 2008). Our empirical evaluations suggest that this mixture is an effective balance between searching for policies that have high cumulative reward and that are admissible. At each SAMPLE step, ARCS traverses a trajectory in $T$, uniform randomly with either heuristic sampling or random sampling. Traversed nodes are selected for Bellman backup in the BACKUP step.
@@
For heuristic sampling, We use the same sampling strategy and sampling termination condition as SARSOP (Kurniawati, Hsu, and Lee 2008), but any sampling technique can be used. It works by choosing actions with the highest $\overline{Q}_{R}$, and observations that makes the largest contribution to the gap at the root of $T$, and sampling terminates based on a combination of selective deep sampling and a gap termination criterion. We refer the reader to (Kurniawati, Hsu, and Lee 2008) for details. With random sampling, actions and observations are chosen at random while traversing the tree until a new node is reached and added to the tree.
Backup
@@
The BACKUP operation at node $v$ updates the values of $k,\overline{Q}_{R},\underline{Q}_{R},\overline{Q}_{C},\underline{Q}_{C},\overline{V}_{R},\underline{V}_{R},\overline{V}_{C}$ and $\underline{V}_{C}$ by back-propagating the information of the children of $v$ back to $v$. First, the values of $\overline{Q}_{R},\underline{Q}_{R},\overline{Q}_{C}$ and $\underline{Q}_{C}$ are computed for each action using Eq. (3) and Eq. (5) for rewards and costs, respectively. Then, an RC-POMDP Bellman Backup Eq. (17) is used to update $\overline{V}_{R},\underline{V}_{R},\overline{V}_{C},\underline{V}_{C}$, by computing the best Q reward-value (and Q cost-value) among actions that satisfy the constraint Eq. (11). The action selected to update $\underline{V}_{R}$ is used to update $k$ by back-propagating the minimum $k$ of all successor nodes. If no actions are feasible, all current policies from that node are inadmissible. In that case, we update the reward- and cost-values using the action with the minimum $Q$-cost value, and set $k=0$.
Pruning
@@
To keep the size of $T$ small and improve tractability, we prune nodes and node-actions that are suboptimal, using the following criteria. First, for each node $v\in B_{sam}$, if $v.\underline{V}_{C}>v.d$, no admissible policies exist from $v$, so $v$ and its subtree are pruned. Next, we prune actions as follows. Let $k(v,a)$ be the admissible horizon guarantees of the successor nodes from taking action $a$ at node $v$. Between two actions $a$ and $a^{\prime}$, if $k(v,a^{\prime})=\infty,v.\overline{Q}_{R}(a)<v.\underline{Q}_{R}(a^{\prime})$, we prune the node-action $(v,a)$ (disallow taking action $a$ at node $v$), since action $a$ can never be taken by the optimal admissible policy. Next, if all node-actions $(v,a)$ are pruned, $v$ is also pruned since no actions are admissible from $v$. Finally, the node-action $(v,a)$ is pruned if any successor node from taking $a$ at $v$ are pruned, since that means $(v,a)$ is also inadmissible. Nodes and node-actions that are pruned are not chosen during action and observation selection during SAMPLE and BACKUP.
@@
ARCS terminates when two conditions are met, (i) when it finds an admissible policy, i.e., $v_{0}.k=\infty$, which is when all leaf nodes $v_{leaf}$ reachable under a policy satisfy either Eq. (19) or $v_{leaf}.V_{C}=0$, and (ii) when it finds an $\epsilon$-optimal policy, i.e., when the gap criterion at the root is satisfied, that is when $v_{0}.\overline{V}_{R}-v_{0}.\underline{V}_{R}\leq\epsilon$.
@@ 5.1 Algorithm Analysis
Here, we analyze the theoretical properties of ARCS. Proofs are provided in the supplementary material.
@@
Given an RC-POMDP with admissibility constraint $k=\infty$, let $T$ be the computed policy tree after some iterations of ARCS. Let $V_{R}^{*}$ and $V_{C}^{*}$ be the reward-value and cost-value of an optimal admissible policy. At every node $v$ with $\bar{b}=(v.b,v.d)$ and admissible horizon guarantee $v.k=\infty$, it holds that:
$v.\underline{V}_{R}\leq V_{R}^{*}(\bar{b})\leq v.\overline{V}_{R}$    and    $v.\underline{V}_{C}\leq V_{C}^{*}(\bar{b})\leq v.\overline{V}_{C}.$
@@
Given an RC-POMDP with admissibility constraint $k=\infty$ and $\epsilon$, if ARCS terminates with a solution, the policy is both an admissible and $\epsilon$-optimal.
ARCS is not complete; It may not terminate for some RC-POMDP problems, due to the conservative computation of admissible horizon and needing to grow an infinitely deep tree to find admissible policies for some problems. However, ARCS can find $\epsilon$-optimal admissible policies for many problems, such as the problems in our evaluation. These are problems where a a finite depth tree is sufficient to compute admissible policies even with the conservatism. We leave the analysis of such classes of RC-POMDPs to future work.
@@ 6 Experimental Evaluation
To the best of our knowledge, this work is the first to propose and solve RC-POMDPs, and thus there are no existing algorithms to compare to directly. The purpose of our evaluation is to (i) empirically compare the behavior of policies computed for RC-POMDPs with those computed for C-POMDPs, and (ii) evaluate the performance of our proposed algorithm for RC-POMDPs. To this end, we consider the following algorithms for our evaluation:
  • CGCP (Walraven and Spaan 2018): State-of-the-art algorithm that computes optimal policies for C-POMDPs.
@@
  • CPBVI (Kim et al. 2011): Approximate dynamic programming that uses admissible cost as a heuristic.
  • CPBVI-D: CPBVI computes stochastic policies. We modify CPBVI to compute deterministic policies to evaluate its efficacy for RC-POMDPs.
  • ARCS (Ours).
We consider the following test environments: (i) CE: Simplified counterexample in Figure 1, (ii) C-Tiger: A constrained version of the Tiger POMDP Problem (Kaelbling, Littman, and Cassandra 1998), with cost of $1$ for the “listen” action, (iii) CRS: A constrained version of the RockSample problem (Smith and Simmons 2004) as defined in (Lee et al. 2018) with varying sizes and number of rocks, and (iv) Tunnels: A scaled version of Example 1, shown in Fig. 2.
@@
CE CGCP $0.51$ $\color[rgb]{0,0,.545}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.545}\pgfsys@color@rgb@stroke{0}{0}{.545}\pgfsys@color@rgb@fill{0}{0}{.545}\mathbf{12.00}$ $5.19$
@@
Ours $\mathbf{0.00}$ $\color[rgb]{0,.392,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,.392,0}\pgfsys@color@rgb@stroke{0}{.392}{0}\pgfsys@color@rgb@fill{0}{.392}{0}\mathbf{10.00}$ $5.00$
C-Tiger CGCP $0.75$ $\mathbf{-1.69}$ $3.00$
CGCP-CL $0.14$ $\color[rgb]{0,0,.545}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.545}\pgfsys@color@rgb@stroke{0}{0}{.545}\pgfsys@color@rgb@fill{0}{0}{.545}\mathbf{-2.98}$ $2.93$
@@
Ours (CL) $\mathbf{0.00}$ $\color[rgb]{0,.392,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,.392,0}\pgfsys@color@rgb@stroke{0}{.392}{0}\pgfsys@color@rgb@fill{0}{.392}{0}\mathbf{-5.75}$ $2.98$
CRS(4,4) CGCP $0.51$ $\color[rgb]{0,0,.545}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.545}\pgfsys@color@rgb@stroke{0}{0}{.545}\pgfsys@color@rgb@fill{0}{0}{.545}\mathbf{10.43}$ $0.51$
@@
Ours $\mathbf{0.00}$ $\color[rgb]{0,.392,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,.392,0}\pgfsys@color@rgb@stroke{0}{.392}{0}\pgfsys@color@rgb@fill{0}{.392}{0}\mathbf{6.52}$ $0.52$
CRS(5,7) CGCP $0.41$ $\color[rgb]{0,0,.545}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.545}\pgfsys@color@rgb@stroke{0}{0}{.545}\pgfsys@color@rgb@fill{0}{0}{.545}\mathbf{11.98}$ $1.00$
@@
Ours $\mathbf{0.00}$ $\color[rgb]{0,.392,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,.392,0}\pgfsys@color@rgb@stroke{0}{.392}{0}\pgfsys@color@rgb@fill{0}{.392}{0}\mathbf{11.77}$ $0.95$
@@
($\hat{c}=1$) CPBVI $0.90$ $\color[rgb]{0,0,.545}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.545}\pgfsys@color@rgb@stroke{0}{0}{.545}\pgfsys@color@rgb@fill{0}{0}{.545}\mathbf{1.92}$ $1.62$
CPBVI-D $0.89$ $\color[rgb]{0,0,.545}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,.545}\pgfsys@color@rgb@stroke{0}{0}{.545}\pgfsys@color@rgb@fill{0}{0}{.545}\mathbf{1.92}$ $1.57$
Ours $\mathbf{0.00}$ $\color[rgb]{0,.392,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,.392,0}\pgfsys@color@rgb@stroke{0}{.392}{0}\pgfsys@color@rgb@fill{0}{.392}{0}\mathbf{1.03}$ $0.44$
Table 1: Comparison of ARCS to state-of-the-art offline C-POMDP algorithms. We report the mean for each metric. We bold the best violation rates in black, the highest reward with violation rate greater than 00 in blue, and the highest reward with violation rate less than 00 in green. Standard error of the mean, environment parameters and experiments with different problem parameters can be found in the appendix.
We implemented each algorithm in Julia under the POMDPs.jl framework (Egorov et al. 2017), and all experiments were conducted single threaded on two nominally $2.2$ GHz Intel Xeon CPUs with $48$ cores and $128$ GB RAM. For all algorithms except CGCP-CL, solve time is limited to $300$ seconds and online action selection to $0.05$ seconds. For CGCP-CL, $300$ seconds was given for each action. We simulate each policy $1000$ times, except for CGCP-CL, which is simulated $100$ times due to the time taken for re-computation of the policy at each time step, and report the average discounted cumulative reward and cost, and constraint violation rate in Table 1. The constraint violation rate is the fraction of trials in which the history dependent admissible cost $d(h_{t})$ becomes negative, which implies that Eq. (13) is violated.
@@
Figure 2: Tunnels Problem. The initial probability of rock existence (red regions) is $0.8,0.4$ and 00 for tunnels $A,B$ and $C$, respectively. There is a cost of $1$ for rock traversal and cost of $0.5$ for backtracking. Trajectories from CGCP (blue) and ARCS (green) are displayed, with opacity approximately proportional to frequency of trajectories.
@@
In all environments, ARCS found admissible policies, i.e., $k=\infty$. On the other hand, CGCP, CPBVI and CPBVI-D could only guarantee an admissible horizon of $k=1$, since the constraint of the C-POMDP formulation is only at the initial time step. CGCP-CL may have a closed-loop admissible horizon greater than 1, but does not provide any guarantees, as indicated in the violation rate.
@@
The benchmarking results show that the policies computed for ARCS generally achieve competitive cumulative reward to policies computed for C-POMDP, without any constraint violations and thus no pathological behavior. ARCS also performs better in all metrics than CPBVI and CPBVI-D, both of which could not search the problem space sufficiently to find good solutions in large RC-POMDPs.
@@
Although the C-POMDP policies generally satisfy the expected cost constraints of Eq. 6, the prevalence of high violation rates of C-POMDP policies across all environments strongly suggests that the manifestation of the stochastic self-destruction for C-POMDP policies is not a rare or exceptional phenomenon, but an inherent aspect of the C-POMDP formulation. This behavior is illustrated in the Tunnels problem, shown in Figure 2. CGCP (in blue) decides to traverse tunnel $A$ $51\%$ of the time even when it observes that $A$ is rocky, and traverses tunnel $B$ $49\%$ of the time. On the other hand, ARCS never traverses tunnel $A$, since such a policy is inadmissible. Instead, it traverses tunnels $B$ and $C$ depending on the observation of rocks in tunnel $B$, to maximize rewards while remaining admissible.
@@
Finally, the closed-loop inconsistency of C-POMDP policies is evident when comparing the open loop CGCP with the closed loop CGCP-CL. In all cases, the cumulative reward is decreased when going from CGCP to CGCP-CL, sometimes drastically. The violation rate also decreases, but not to 00, suggesting that planning with C-POMDPs instead of RC-POMDPs can lead to irrecoverable behavior that cannot be addressed by closed-loop re-planning. As seen in CE and both CRS, closed-loop CGCP-CL also gets lower reward than ARCS while still having constraint violations. This implies that, even for closed-loop planning, RC-POMDP is more advantageous than C-POMDP.
7 Conclusion

### v3

@@ 5 Dynamic Programming for RC-POMDPs
where $\alpha_{c}$ refers to a cost $\alpha$-vector. This can be solved efficiently with the maximin LP Williams [1990]:
$\max_{z,b}\,z\;\;\text{ s.t. }\;\;\;\alpha_{c}^{T}b\geq z,\;\;(\alpha_{r},\alpha_{c})\in\Gamma,\;\;b\in\Delta(S).$ (22)
@@
Sampling   ARCS uses a mixture of random sampling Spaan and Vlassis [2005] and heuristic search (SARSOP) Kurniawati et al. [2008]. Our empirical evaluations suggest that this approach is an effective balance between finding policies with high cumulative reward and that are admissible. At each SAMPLE step, ARCS expands the search space from the root of $T$, with either heuristic sampling or random sampling. For heuristic sampling, we use the same sampling strategy and sampling termination condition as SARSOP. It works by choosing actions with the highest $\overline{Q}_{R}$, and observations that have the largest contribution to the gap at the root of $T$, and sampling terminates based on a combination of selective deep sampling and a gap termination criterion. With random sampling, actions and observations are chosen randomly while traversing the tree until a new node is reached and added to the tree. Sampled points are chosen for BACKUP.
@@
Backup   The BACKUP operation at node $v$ updates the values in the node by back-propagating the information of the children of $v$ back to $v$. First, the values of $\overline{Q}_{R},\underline{Q}_{R}$ and $\overline{Q}_{C}$ are computed for each action using Eq. (3) and Eq. (5) for rewards and costs, respectively. Then, an RC-POMDP backup Eq. (15) is used to update $\overline{V}_{R},\underline{V}_{R},\underline{V}_{C},\overline{V}_{C}$. The action selected to update $\underline{V}_{R}$ is used to update $k$ by back-propagating the minimum $k$ of all children. If no actions are feasible, all current policies from that node are inadmissible, and we update the reward- and cost-values using the action with the minimum $Q$-cost value, and set $k=0$.
@@
Pruning   To keep the size of $T$ small and improve tractability, we prune nodes and node-actions that are suboptimal, using the following criteria. First, for each node $v\in B_{sam}$, if $v.\underline{V}_{C}>v.d$, no admissible policies exist from $v$, so $v$ and its subtree are pruned. Next, we prune actions as follows. Let $k(v,a)$ be the admissible horizon guarantees of the successor nodes from taking action $a$ at node $v$. Between two actions $a$ and $a^{\prime}$, if $k(v,a^{\prime})=\infty\text{ and }v.\overline{Q}_{R}(a)<v.\underline{Q}_{R}(a^{\prime})$, we prune the node-action $(v,a)$ (disallow taking action $a$ at node $v$), since action $a$ can never be taken by the optimal admissible policy. Next, if all node-actions $(v,a)$ are pruned, $v$ is also pruned. Finally, the node-action $(v,a)$ is pruned if any successor node from taking $a$ at $v$ is pruned. Nodes and node-actions that are pruned are not chosen during action and observation selection during SAMPLE and BACKUP.
@@
ARCS terminates when two conditions are met, (i) when it finds an admissible policy, i.e., $v_{0}.k=\infty$, which is when all leaf nodes $v_{leaf}$ reachable under a policy satisfy Eq. (20), and (ii) when it finds an $\epsilon$-optimal policy, i.e., when the gap criterion at the root is satisfied, that is when $v_{0}.\overline{V}_{R}-v_{0}.\underline{V}_{R}\leq\epsilon$.
@@ 5.1 Algorithm Analysis
Here, we analyze the theoretical properties of ARCS.
@@
Given an RC-POMDP with admissibility constraint $k=\infty$, let $T$ be the policy tree after some iterations of ARCS. Let $V_{R}^{*}$ be the reward-value of an optimal admissible policy. At every node $v$ with $\bar{b}=(b,d)$ and admissible horizon guarantee $v.k=\infty$, it holds that:
$v.\underline{V}_{R}\leq V_{R}^{*}(\bar{b})\leq v.\overline{V}_{R}$    and    $v.\overline{V}_{C}\leq d.$
@@
Given an RC-POMDP with admissibility constraint $k=\infty$ and $\epsilon$, if ARCS terminates with a solution, the policy is admissible and $\epsilon$-optimal.
ARCS is not complete. It may not terminate, due to conservative computation of admissible horizon and needing to search infinitely deep to find admissible policies for some problems. However, ARCS can find $\epsilon$-optimal admissible policies for many problems, such as the ones in our evaluation. These are problems where a finite depth is sufficient to compute admissibility even with conservatism. We leave the analysis of such classes of RC-POMDPs to future work.
@@ 6 Experimental Evaluation
To the best of our knowledge, this work is the first to propose and solve RC-POMDPs, and thus there are no existing algorithms to compare to directly. The purpose of our evaluation is to (i) empirically compare the behavior of policies computed for RC-POMDPs with those computed for C-POMDPs, and (ii) evaluate the performance of our proposed algorithm for RC-POMDPs. To this end, we consider the following offline algorithms to compare against our ARCS [^3] Our code is open sourced at https://github.com/CU-ADCL/RC-PBVI.jl:
  • CGCP Walraven and Spaan [2018]: Algorithm that computes near-optimal policies for C-POMDPs using a primal-dual approach.
@@
  • Exp-Gradient Kalagarla et al. [2022]: Algorithm that computes mixed policies for C-POMDPs using a no-regret learning approach with a primal-dual approach using an Exponentiated Gradient method.
  • CPBVI Kim et al. [2011]: Approximate dynamic programming that uses admissible cost as a heuristic.
  • CPBVI-D: We modify CPBVI to compute deterministic policies to evaluate its efficacy for RC-POMDPs.
Since the purpose of our comparison between RC-POMDPs and C-POMDPs is mainly with regard to constraints, we do not compare to online C-POMDP algorithms such as CC-POMCP Lee et al. [2018] which can handle larger problems but do not have anytime guarantees on constraint satisfaction.
We consider the following environments: (i) CE: Counterexample in Figure 1, (ii) C-Tiger: A Constrained version of Tiger POMDP Kaelbling et al. [1998], (iii) CRS: Constrained RockSample Lee et al. [2018], and (iv) Tunnels: A scaled version of Example 1, shown in Figure 2. Details on each problem, experimental setup, and algorithm implementation are in the Appendix. For all algorithms except CGCP-CL, solve time is limited to $300$ seconds and online action selection to $0.05$ seconds. For CGCP-CL, $300$ seconds was given to re-compute each action. We report the mean discounted cumulative reward and cost, and constraint violation rate in Table 1. The constraint violation rate is the fraction of trials in which $d(h_{t})$ becomes negative, which means Eq. (11) is violated.
Table 1: Results for benchmarks. We report the mean for each metric. We bold the best violation rates in black, the highest reward with violation rate greater than 00 in blue, and the highest reward with 00 violation rate in green. Standard error of the mean, and problem parameters can be found in the appendix.
@@
CE CGCP $0.51$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbf{12.00}$ $5.19$
@@
Ours $\mathbf{0.00}$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbf{10.00}$ $5.00$
C-Tiger CGCP $0.75$ $-1.69$ $3.00$
CGCP-CL $0.14$ $-2.98$ $2.93$
Exp-Gradient 1.0 $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbf{1.81}$ 3.22
@@
Ours $\mathbf{0.00}$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbf{-5.75}$ $2.98$
CRS(4,4) CGCP $0.51$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbf{10.43}$ $0.51$
@@
Ours $\mathbf{0.00}$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbf{6.52}$ $0.52$
CRS(5,7) CGCP $0.41$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbf{11.98}$ $1.00$
@@
Ours $\mathbf{0.00}$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbf{11.77}$ $0.95$
CRS(7,8) CGCP $0.36$ $10.78$ $0.945$
CL-CGCP $0.20$ $\mathbf{11.17}$ $0.931$
($\hat{c}=1$) EXP-Gradient 0.32 10.03 1.15
CPBVI $\mathbf{0.00}$ $0.0$ $0.0$
CPBVI-D $\mathbf{0.00}$ $0.0$ $0.0$
Ours $\mathbf{0.00}$ $\mathbf{6.61}$ $0.960$
@@
($\hat{c}=1$) Exp-Gradient 0.48 1.35 0.82
CPBVI $0.90$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbf{1.92}$ $1.62$
CPBVI-D $0.89$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbf{1.92}$ $1.57$
Ours $\mathbf{0.00}$ $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbf{1.03}$ $0.44$
@@
Figure 2: Tunnels. There is a cost of $1$ for rock traversal (red regions) and $0.5$ for backtracking. Trajectories from CGCP (blue) and ARCS (green) are displayed, with opacity approximately proportional to frequency of trajectories.
@@
In all environments, ARCS found admissible policies ($k=\infty$). In contrast, CGCP, Exp-Gradient, CPBVI and CPBVI-D only guarantee an admissible horizon of $k=1$, since the C-POMDP constraint is only at the initial belief. CGCP-CL may have a closed-loop admissible horizon greater than 1, but does not provide guarantees, as indicated in the violation rate.
@@
The benchmarking results show that the policies computed for ARCS generally achieve competitive cumulative reward to policies computed for C-POMDP, without any constraint violations and thus no pathological behavior. ARCS also generally performs better in all metrics than CPBVI and CPBVI-D, both of which could not search the problem space sufficiently to find good solutions in large RC-POMDPs.
@@
Although the C-POMDP policies generally satisfy the C-POMDP expected cost constraints, the prevalence of high violation rates of C-POMDP policies across the environments strongly suggests that the manifestation of the stochastic self-destruction in C-POMDPs is not an exceptional phenomenon, but intrinsic to the C-POMDP problem formulation. This behavior is illustrated in the Tunnels problem, shown in Figure 2. CGCP (in blue) decides to traverse tunnel $A$ $51\%$ of the time even when it observes that $A$ is rocky, and traverses tunnel $B$ $49\%$ of the time. In contrast, ARCS never traverses tunnel $A$, since such a policy is inadmissible. Instead, it traverses $B$ or $C$ depending on observation of rocks in tunnel $B$, to maximize rewards while remaining admissible.
@@
Finally, the closed-loop inconsistency of C-POMDP policies is evident when comparing open loop CGCP with closed loop CGCP-CL. In most cases (all except CRS(7,8)), the cumulative reward is decreased when going from CGCP to CGCP-CL, sometimes drastically. The violation rate also decreases, but not to 00, suggesting that planning with C-POMDPs instead of RC-POMDPs can lead to myopic behavior that cannot be addressed by re-planning. As seen in CE and both CRS, CGCP-CL attains lower reward than ARCS while still having constraint violations. Therefore, even for closed-loop planning, RC-POMDP can be more advantageous than C-POMDP.
Unconstrained POMDP problems
Table 2: Results for computed policy under-approximation (lower bound for reward values and upper bounds for cost values), best reward values in bold. SARSOP only considers reward value as an unconstrained POMDP algorithm.
Env. Algorithm Reward Cost
CE SARSOP (POMDP) $\mathbf{12.0}$
CGCP (C-POMDP) $\mathbf{12.0}$ $5.0$
$300s$ Ours (RC-POMDP) $\mathbf{12.0}$ $5.0$
C-Tiger SARSOP (POMDP) $\mathbf{1.93}$
CGCP (C-POMDP) $1.90$ 3.2
$300s$ Ours (RC-POMDP) -1.4 3.2
Tunnels SARSOP (POMDP) $\mathbf{1.92}$
CGCP (C-POMDP) $\mathbf{1.92}$ 1.6
$300s$ Ours (RC-POMDP) $\mathbf{1.92}$ 1.6
CRS(4,4) SARSOP (POMDP) $\mathbf{16.9}$
CGCP (C-POMDP) $\mathbf{16.9}$ 2.4
$300s$ Ours (RC-POMDP) $\mathbf{16.9}$ 2.2
CRS(5,7) SARSOP (POMDP) $\mathbf{23.9}$
CGCP (C-POMDP) $14.8$ $3.6$
$300s$ Ours (RC-POMDP) $14.9$ $2.1$
CRS(5,7) SARSOP (POMDP) $\mathbf{24.0}$
CGCP (C-POMDP) $24.0$ $4.5$
$1000s$ Ours (RC-POMDP) $15.3$ $2.2$
Next, we evaluate how well the RC-POMDP framework and our proposed algorithm performs for problems that have reduced constraints, so as to trivially become an unconstrained POMDP. We evaluate ARCS (RC-POMDP), CGCP (C-POMDP algorithm) and SARSOP (unconstrained POMDP algorithm) for the same benchmark problems with very high constraint thresholds $\hat{c}=1000$.
For these problems, all policies are admissible, and ARCS is guaranteed to asymptotically converge to the optimal solution. However, since ARCS needs to keep track of admissible cost values, we utilize a policy tree representation. This representation is less efficient than the $\alpha$-vector policy representation used in SARSOP and CGCP, which allow value improvements at a belief state to directly improve values at other belief states.
Table 2 reports the lower bound reward and upper bound costs computed by each algorithm with a time limit of $300s$. As seen in Table 2, our algorithm performs similar to CGCP and the unconstrained POMDP algorithm SARSOP for most smaller problems. The C-Tiger problem benefits greatly from the $\alpha$-vector representation, since the optimal policy repeatedly cycles among a small set of belief states (which our algorithm considers different augmented belief-admissible cost states). For slightly larger problems (CRS(5,7)), the efficient $\alpha$-vector representation and other heuristics of SARSOP (which CGCP takes advantage of, since it repeatedly calls SARSOP) enables much faster convergence than the policy tree-based method of our approach. Nonetheless, as time is increased, our algorithm slowly improves its values.
Overall, ARCS exhibits competitive performance in problems with reduced or no constraints, albeit with less scalability. However, the main advantage of the RC-POMDP formulation is in the careful treatment of constraints to mitigate the pathological behaviors of the C-POMDP formulation, making RC-POMDPs particularly valuable for problems involving nontrivial constraints.
7 Conclusion and Future Work